# Smooth surface reconstruction of earthquake faults from distributed moment-potency-tensor solutionsC

### v2

@@ 3 Disaster Prevention Research Institute
The earthquake fault as observed by seismic motion primarily manifests as a surface of displacement discontinuity within a linear elastic continuum. The displacement discontinuity and the surface normal vector ($n$-vector) of this idealized earthquake source are measured by the tensor of potency, which is seismic moment normalized by stiffness. We exploit this theoretical relation to formulate an inverse problem of reconstructing a smooth, three-dimensional fault surface from an areal density field of the potency tensor. In this problem, the surface is represented by an elevation field that parametrizes the vertical variation of the surface relative to a reference, and the nodal planes of a given potency-density-tensor field describe the $n$-vector field. The remaining subject is the $n$-vector-to-elevation transform, the operation inverse to defining the $n$-vector field on a given surface. Whereas this transform is a well-posed one-to-one mapping in two dimensions where the $n$-vector has one degree of freedom, the transform becomes overdetermined in three dimensions because the $n$-vector has two degrees of freedom while the scalar elevation has only one, generally admitting no solution. This overdetermination originates from a reduction in degrees of freedom from six to five upon modeling the source as a displacement discontinuity rather than general potency density, namely inelastic strain. The sixth degree of freedom unmodeled by displacement discontinuities and $n$-vectors manifests as a localized violation of the determinant-free constraint in point potency sources; however, in areal sources of potency density, it appears as a conflict with the global consistency of the $n$-vector field. Recognizing that this conflict derives from the capacity of the potency-density-tensor field to describe inelastic strain source incompatible with displacement discontinuity on a surface, we explicitly introduce an a priori constraint to define the fault surface as the smooth surface that best approximates the surface distribution of inelastic strain by displacement discontinuity. We derive an analytical solution for the surface reconstruction thus formulated and demonstrate its ability to reproduce smooth three-dimensional surfaces from synthetic noisy $n$-vector fields. Lastly, we integrate the derived formula into the potency density tensor inversion and validate it in an application to the 2013 Balochistan earthquake. The estimated fault geometry agrees better with the observed fault trace than that of the previously proposed quasi-two-dimensional surface reconstruction, highlighting the importance of accounting for three-dimensional fault geometry.
@@ 1 Preliminaries
A series of developments are based on slip inversion (Chinnery, 1964), which estimates an earthquake source as a displacement discontinuity on a boundary (or its gradient, dislocation; Matsu’ura, 1977a, b). Slip inversion typically confines the model-parameter space to shear slip (Chinnery, 1964), while variants exist that estimate tensile faults as well (Massonnet and Feigl, 1998). Either form of analysis could be termed slip-and-opening inversion or displacement-discontinuity inversion; however, as no definitive nomenclature has been established, this paper collectively refers to all inverse problems that model the earthquake source as a displacement discontinuity as ‘slip inversion’ in a broad sense. Within the framework of slip inversion, slip (and opening) distributions on faults are model parameters, and fault geometry is an additional set of parameters (Savage and Hastie, 1966; Matsu’ura and Hasegawa, 1987; Fukahata and Wright, 2008; Ragon et al., 2018) that determine the structure of the inference model, associated with hyperparameters in statistics (Matsu’ura, 1977a). The simplest, highly successful example along this line is configuration estimation using a rectangular fault, which infers the center of masses, angles, and lengths of a fault, besides the direction of the displacement discontinuity (Savage and Hastie, 1966; Matsu’ura and Hasegawa, 1987). Detailed slip patterns can be estimated from a distributed slip inversion on a prescribed fault surface (Yabuki and Matsu’ura, 1992). When nonplanar geometry is inferred in slip inversion, fault subdivisions with a finite number of planar subfaults have been widely utilized (Matsu’ura, 1977a, b; Loveless and Meade, 2010; Avouac et al., 2014). This approach, however, suffers from the artificial stress singularity arising at subfault edges (Romanet et al., 2020, 2024; Mallick and Meade, 2025). A primary objective of fault geometry inference is the estimation of the macroscopic fault geometry, which is coarse-grained as a smooth surface consistent with linear continuum mechanics (Matsu’ura et al., 2019; Dutta et al., 2021).
@@
In slip inversion, where the estimation of displacement discontinuity is a linear problem, the estimation of fault geometry alters the problem into a nonlinear inference since the kernel function (Green’s function) to relate slip and observed data depends on the geometry (Yabuki and Matsu’ura, 1992). This nonlinearity contrasts with the CMT analysis, where far-field conditions enabling the point-source approximation ensure the positional insensitivity of Green’s functions, allowing for stable perturbative updates from a coarse initial solution (Dziewonski et al., 1981). The radiation pattern of Green’s function amplitudes, which conveys source geometry information to the far-field, is determined by the potency representing the slip direction and the fault normal, rendering the CMT analysis an almost linearly-solvable inversion. In turn, this quasilinearity holds without the point-source approximation if the source is formulated as a density field of such potency. This realization paved the way for the potency density tensor inversion (PDTI, Kikuchi and Kanamori, 1991; Kasahara and Yagi, 2010; Shimizu et al., 2020), where the earthquake fault is expressed as a potency-density-tensor field distributed over a domain of negligible volume, a pseudo boundary.
@@
In the PDTI, the model parameters correspond to the areal density field of the potency tensor set along a pseudo boundary. The displacement discontinuity and the fault normal are represented by the nodal planes of the estimated potency density tensors that smoothly rotate over space and time, where those two are equivalent information. By permitting potency incompatible with the normal of the pseudo boundary placed as a first-order approximation of the actual fault, the PDTI extracts geometry information consistent with observed seismic motions. Because the seismic moment density represents the deviation of stress from elasticity (stress glut, Backus and Mulcahy, 1976a, b), the seismic moment normalized by the stiffness, the potency density describes the inelastic strain (Kostrov, 1974). Consequently, the potency density tensor possesses six degrees of freedom as long as it is treated as a representation of inelastic strain within a finite volume, although it is essentially equivalent to the tensor product of a displacement discontinuity (slip and opening) and a fault normal in the limit of infinitesimal volumes; recall that the strain tensor has six degrees of freedom whereas the set of fault normal and displacement discontinuity has five degrees of freedom due to the normalization condition of the $n$-vector. Although counterintuitive, the inelastic strain source, the potency density, is intrinsically higher-dimensional than displacement discontinuity. Therefore, the PDTI can express multiple earthquake faults, such as a conjugate fault system and a set of disconnected faults, on a single model plane (Yamashita et al., 2021; Okuwaki and Fan, 2022), enabling the modeling of proximal faults even when the inter-surface distance falls below the observation limit (Volkov and Triki, 2025) of discernibility.
@@
The pioneering research of Kikuchi and Kanamori (1991) first infers a seismic source as inelastic strain, namely the potency density, distributed over a pseudo boundary instead of as displacement discontinuity across a volume-less interface. Nonetheless, their original PDTI is almost indistinguishable from planar fault subdivision and similar to multiple point source approaches (Tsai et al., 2005; Yue and Lay, 2020) because the prior constraint they employ is only the double-couple condition imposed as a traceless and determinant-free constraint, where the determinant-free constraint functions as the point-source displacement-discontinuity condition while the traceless condition eliminates the volumetric deformation. A major advantage of the PDTI stems from the affinity for incorporating a prior constraint upon the fault geometry, expressible as the hyperprior onto geometrical hyperparameters in the context of slip inversion, which is, to our knowledge, first introduced by Kasahara and Yagi (2010). The prior for the potency operates on both the displacement discontinuity and the $n$-vector field, and thus, smoothly varying macroscopic rotations of focal mechanisms are extracted by simply using a smoothness constraint on the model parameter in the PDTI. Kasahara and Yagi (2010) have further developed a stable scheme by incorporating the propagation of Green’s function errors (Yagi and Fukahata, 2010, 2011) typified by structural uncertainties, which induce apparent rotations of radiation patterns for near-vertical take-off angles (Ford et al., 2012; Chiang et al., 2014) and travel-time perturbations (Hu et al., 2023, 2025). Whereas the original intention of Kasahara and Yagi (2010) was rather to extract spatial variations in slip vectors and focal mechanisms themselves, Shimizu et al. (2020) point out that spatial patterns of the potency focal mechanism also document fault geometry as fault normal vector fields. The PDTI untangles source complexities in the perspective of fault geometry thereafter (Okuwaki et al., 2020; Tadapansawut et al., 2021; Fan et al., 2022; Ohara et al., 2023; Yagi et al., 2024; Inoue et al., 2025).
@@ 2.1 Observation equations
In the PDTI, $D$ is defined via eq. (1) as a source representation that generally possesses six degrees of freedom (DOFs) relative to observations, implying that eq. (2) is not strictly satisfied. A symmetric matrix composed of two vectors, as in eq. (2), has a rank of at most two and is subject to a determinant-free constraint, meaning that a displacement-discontinuity source cannot exceed five DOFs. Indeed, the combined DOFs of a displacement discontinuity and an $n$-vector amount to five. Equation (2) restricts the opening mode to the fault-normal direction by imposing vanishing tangential-tangential components. Although a traceless condition is typically imposed to exclude tensile components when assuming pure shear slip, the tensile component by default serves to resolve the degeneracy between the fault normal and the displacement discontinuity within the moment/potency tensor.
Equation (2) is derived as an integrability condition, namely the compatibility condition in continuum mechanics, for the surface distribution of inelastic strain (Appendix A). There is thus no physical requirement for it to be satisfied as long as the source is treated as inelastic strain within a finite volume. Therefore, eq. (1) should not be read as a mere restatement of the representation theorem for displacement discontinuities. It is instead a representation theorem for finite-volume sources where thickness is neglected. As demonstrated in the following section, the primary objective in geometry reconstruction from the potency is to extract the surface normal $n$ (eq. 3), defined in the displacement-discontinuity limit, from a source that is not strictly an ideal surface. For a point source, this task reduces to ignoring the eigenvalue with the smallest absolute magnitude to satisfy the determinant-free condition. For an areal source, it requires addressing the global constraints of eq. (3), which act through infinitesimal variations along the surface.
@@ 2.2 Decoupling surface reconstruction from potency estimation
The PDTI evaluates the potency density $D$ and the associated $n$-vector field through the well-known equations (1) and (2), and the $n$-vector field is translated into $\Gamma$ through eq. (3). Here is a problematic circular dependency between $D$ and $\Gamma$: the domain on which the potency density $D$ is defined is the fault surface $\Gamma$ (eq. 1), which is yet to be evaluated by $D$ (eq. 3). To infer $D$ prior to reconstructing $\Gamma$, the representation theorem (eq. 1) must be projected onto a fixed surface, here termed the reference surface denoted by $\Gamma^{\circ}$ (Fig. 1).
@@ 3 Formulation of inverse mapping from potency density to fault shape
Equation (7) states that the cotangent of the $n$-vector corresponds to the downward gradient of a curve in a two-dimensional space, where the slip direction perpendicular to the $n$-vector is exactly the fault orientation. Accordingly, Shimizu et al. (2021) averaged the potency along the short axis of the reference surface and smoothly connected the resulting $n$-vector along the long axis to extract the two-dimensional features of a uniaxially bended fault (e.g., along-strike strike angle variations and along-dip dip angle variations). The same procedure adhering to eq. (6) can also reproduce quasi-two-dimensional shapes including uniaxial twists such as along-strike dip variations and along-dip strike variations (§5.3). This study regards such surface reconstruction fully governed by eq. (6), including quasi-two-dimensional surface reconstruction, as falling within the scope of the previous method by Shimizu et al. (2021).
@@
Then, eq. (13) requires the limit that places $\bf n$} in the integrable nearest neighbor of non-integrable $\bf n$$_{*}$, that is, the displacement discontinuity that best approximates given inelastic strain. It is termed dislocation limit hereafter, which is, in eq. (14),
@@
When the non-integrable part $(\partial_{x}\partial_{y}-\partial_{y}\partial_{x})z$ expected from $n_{*}$ is small, the vector $n$ obtained from eq. (13) does not deviate significantly from $n_{*}$. Basically, we suppose this regime and proceed by adopting the prior $P(\Gamma|{\bf a})$ written as eq. (14). However, this inherent overconstraint in surface reconstruction, where no smooth surface satisfies the forward model eq. (2) that leads to eq. (6), raises the question of what priors $P(\Gamma|{\bf a})$ are appropriate to relax eq. (2) when the non-integrability of $n_{*}$ is not small. It is important to note that the a priori constraint here is imposed on $\Gamma$ given $\bf a$}, rather than on the inelastic strain $\bf a$} distributed over a given surface $\Gamma$ as in slip inversion. Considering that $\bf a$} represents the $n$-vector, this approach of $\Gamma|{\bf a}$ implies a reversal of the standard definition of the $n$-vector, where $n$ is defined for $\Gamma(z)$. The PDTI is commonly formulated in such a way that determines $\Gamma$ from $\bf a$}, which is inferred by using another prior constraint imposed on $\bf a$} independently of $\Gamma$. $P(\Gamma|{\bf a})$ introduced here was designed to align with this structure. While this approach reverses the standard definition of the $n$-vector, the physical nature of faulting supports this causal order of $\Gamma|{\bf a}$ rather than contradicts it. Before the $n$-vector is disturbed by measurement errors that do not satisfy eq. (6), an earthquake source as a fault surface is an approximation of a fracture zone, reflecting the nature of brittle failure localized within a quasi-face (Matsu’ura et al., 2019), so the exact observance of eq. (2) is unphysical from the outset. Thus, if read as a forward model that precedes the observation, $P(\Gamma|{\bf a})$ for returning a surface from a given $n$-vector field is a description of the ambiguity inherent in determining an earthquake fault surface, that is, an ideal model for a thin sheet of inelastic deformation.
@@
In summary, the potency density that best explains the data retains information that cannot be mapped back onto a displacement discontinuity, a limit of inelastic strain strictly localized on a surface. The resulting surface inevitably violates integrability, which is the incompatibility of the potency density with a displacement-discontinuity representation. Briefly, the best earthquake source is generally not a displacement discontinuity. In conventional point-source representations, the discrepancy between a six-degree-of-freedom moment tensor and a five-degree-of-freedom dislocation is resolved locally by simply discarding the eigenvalue with the smallest absolute magnitude. By contrast, in an areal finite-source representation, this dimensionality gap manifests as a non-local geometric conflict. Physically, when a normal vector $n$ is assigned to each coordinate based on local potency, the resulting field fails to satisfy the gradient-based constraints that couple adjacent points. The reconstructed source thus represents not a smooth, continuous fault, but a collection of infinitesimal fragments that are “torn everywhere”, a fragmented geometry where the dislocation is nowhere integrable (eq. 12). Therefore, reconstructing a meaningful fault geometry requires more than a local approximation, necessitating a global approach to the overdeterminacy inherent in the transition from an areal inelastic strain field to a displacement-discontinuity surface. Equation (11) represents the compatibility condition necessary to project the areal inelastic strain onto displacement discontinuity without loss of information. It ensures the global consistency of the normal vector field, functioning as the areal-source counterpart to the determinant-free condition in point-source analysis.

### v3

@@ 3 Disaster Prevention Research Institute
The earthquake fault as observed by seismic motion primarily manifests as a surface of displacement discontinuity within a linear elastic continuum. The displacement discontinuity and the surface normal vector ($n$-vector) of this idealized earthquake source are measured by the tensor of potency, which is seismic moment normalized by stiffness. We exploit this theoretical relation to formulate an inverse problem of reconstructing a smooth, three-dimensional fault surface from an areal density field of the potency tensor. In this problem, the surface is represented by an elevation field that parametrizes the vertical variation of the surface relative to a reference, and the nodal planes of a given potency-density-tensor field describe the $n$-vector field. The remaining subject is the $n$-vector-to-elevation transform, the operation inverse to defining the $n$-vector field on a given surface. Whereas this transform is a well-posed one-to-one mapping in two dimensions where the $n$-vector has one degree of freedom, the transform becomes overdetermined in three dimensions because the $n$-vector has two degrees of freedom while the scalar elevation has only one, generally admitting no solution. This overdetermination originates from a reduction in degrees of freedom from six to five upon modeling the source as a displacement discontinuity rather than general potency density, namely inelastic strain. The sixth degree of freedom unmodeled by displacement discontinuities and $n$-vectors manifests as a local violation of the determinant-free constraint in point potency sources; however, in areal sources of potency density, it raises a conflict with the global consistency of the $n$-vector field. Recognizing that this conflict derives from the capacity of the potency-density-tensor field to describe inelastic strain source incompatible with displacement discontinuity on a surface, we explicitly introduce an a priori constraint to define the fault surface as the smooth surface that best approximates the surface distribution of inelastic strain by displacement discontinuity. We derive an analytical solution for the surface reconstruction thus formulated and demonstrate its ability to reproduce smooth three-dimensional surfaces from synthetic noisy $n$-vector fields. Lastly, we integrate the derived formula into the potency density tensor inversion and validate it in an application to the 2013 Balochistan earthquake. The estimated fault geometry agrees better with the observed fault trace than that of the previously proposed quasi-two-dimensional surface reconstruction, highlighting the importance of accounting for three-dimensional fault geometry.
@@ 1 Preliminaries
A series of developments are based on slip inversion (Chinnery, 1964), which estimates an earthquake source as a displacement discontinuity on a boundary (or its gradient, dislocation; Matsu’ura, 1977a, b). Slip inversion typically confines the model-parameter space to shear slip (Chinnery, 1964), while variants exist that estimate tensile faults as well (Massonnet and Feigl, 1998). Either form of analysis could be termed slip-and-opening inversion or displacement-discontinuity inversion; however, as no definitive nomenclature has been established, this paper collectively refers to all inverse problems that model the earthquake source as a displacement discontinuity as ‘slip inversion’ in a broad sense. Within the framework of slip inversion, slip (and opening) distributions on faults are model parameters, and fault geometry is an additional set of parameters (Savage and Hastie, 1966; Matsu’ura and Hasegawa, 1987; Fukahata and Wright, 2008; Ragon et al., 2018) that determine the structure of the inference model, associated with hyperparameters in statistics (Matsu’ura, 1977a). The simplest, highly successful example along this line is configuration estimation using a rectangular fault, which infers the center of masses, angles, and lengths of a fault, besides the vector of the displacement discontinuity (Savage and Hastie, 1966; Matsu’ura and Hasegawa, 1987). Detailed slip patterns can be estimated from a distributed slip inversion on a prescribed fault surface (Yabuki and Matsu’ura, 1992). When nonplanar geometry is inferred in slip inversion, fault subdivisions with a finite number of planar subfaults have been widely utilized (Matsu’ura, 1977a, b; Loveless and Meade, 2010; Avouac et al., 2014). This approach, however, suffers from the artificial stress singularity arising at subfault edges (Romanet et al., 2020, 2024; Mallick and Meade, 2025). A primary objective of fault geometry inference is the estimation of the macroscopic fault geometry, which is coarse-grained as a smooth surface consistent with linear continuum mechanics (Matsu’ura et al., 2019; Dutta et al., 2021).
@@
In slip inversion, where the estimation of displacement discontinuity is a linear problem, the estimation of fault geometry alters the problem into a nonlinear inference since the kernel function (Green’s function) to relate slip and observed data depends on the geometry (Yabuki and Matsu’ura, 1992). This nonlinearity contrasts with the CMT analysis, where far-field conditions enabling the point-source approximation ensure the positional insensitivity of Green’s functions, allowing for stable perturbative updates from a coarse initial solution (Dziewonski et al., 1981). The radiation pattern of Green’s function amplitudes, which conveys source geometry information to the far-field, is determined by the potency representing the slip vector and the fault normal, rendering the CMT analysis an almost linearly-solvable inversion. In turn, this quasilinearity holds without the point-source approximation if the source is formulated as a density field of such potency. This realization paved the way for the potency density tensor inversion (PDTI, Kikuchi and Kanamori, 1991; Kasahara and Yagi, 2010; Shimizu et al., 2020), where the earthquake fault is expressed as a potency-density-tensor field distributed over a domain of negligible volume, a pseudo boundary.
@@
In the PDTI, the model parameters correspond to the areal density field of the potency tensor set along a pseudo boundary. The displacement discontinuity and the fault normal are equivalently expressed by the nodal planes of the estimated potency density tensors that smoothly rotate over space and time. By permitting potency incompatible with the normal of the pseudo boundary placed as a first-order approximation of the actual fault, the PDTI extracts geometry information consistent with observed seismic motions. Because the seismic moment density represents the deviation of stress from elasticity (stress glut, Backus and Mulcahy, 1976a, b), the seismic moment density normalized by the stiffness, the potency density describes the inelastic strain (Kostrov, 1974). Consequently, the potency density tensor possesses six degrees of freedom as long as it is treated as a representation of inelastic strain within a finite volume, although it is essentially equivalent to the tensor product of a displacement discontinuity (slip and opening) and a fault normal in the limit of infinitesimal volumes; recall that the strain tensor has six degrees of freedom whereas the set of fault normal and displacement discontinuity has five degrees of freedom due to the normalization condition of the $n$-vector. Although counterintuitive, the inelastic strain source, the potency density, is intrinsically higher-dimensional than displacement discontinuity. Therefore, the PDTI can express multiple earthquake faults, such as a conjugate fault system and a set of disconnected faults, on a single model plane (Yamashita et al., 2021; Okuwaki and Fan, 2022), enabling the modeling of proximal faults even when the inter-surface distance falls below the observation limit (Volkov and Triki, 2025) of discernibility.
@@
The pioneering research of Kikuchi and Kanamori (1991) first inferred a seismic source as inelastic strain, namely the potency density, distributed along a pseudo boundary instead of as displacement discontinuity across a volume-less interface. Nonetheless, their original PDTI is almost indistinguishable from planar fault subdivision and similar to multiple point source approaches (Tsai et al., 2005; Yue and Lay, 2020) because the prior constraint they employ is only the double-couple condition imposed as a traceless and determinant-free constraint, where the determinant-free constraint functions as the point-source displacement-discontinuity condition while the traceless condition eliminates the volumetric deformation. A major advantage of the PDTI stems from the affinity for incorporating a prior constraint upon the fault geometry, expressible as the hyperprior imposed onto geometrical hyperparameters in the framework of slip inversion, which is, to our knowledge, first introduced by Kasahara and Yagi (2010). The prior for the potency operates on both the displacement discontinuity and the $n$-vector field, and thus, smoothly varying macroscopic rotations of focal mechanisms are extracted by simply using a smoothness constraint on the model parameter in the PDTI. Kasahara and Yagi (2010) have further developed a stable scheme by incorporating the propagation of Green’s function errors (Yagi and Fukahata, 2010, 2011) typified by structural uncertainties, which induce apparent rotations of radiation patterns for near-vertical take-off angles (Ford et al., 2012; Chiang et al., 2014) and travel-time perturbations (Hu et al., 2023, 2025). Whereas the original intention of Kasahara and Yagi (2010) was rather to extract spatial variations in slip vectors and focal mechanisms themselves, Shimizu et al. (2020) point out that spatial patterns of the potency focal mechanism also document fault geometry as fault normal vector fields. The PDTI untangles source complexities in the perspective of fault geometry thereafter (Okuwaki et al., 2020; Tadapansawut et al., 2021; Fan et al., 2022; Ohara et al., 2023; Yagi et al., 2024; Inoue et al., 2025).
@@ 2.1 Observation equations
In the PDTI, $D$ is defined via eq. (1) as a source representation that generally possesses six degrees of freedom (DOFs) relative to observations, implying that eq. (2) is not strictly satisfied. A symmetric matrix composed of two vectors, as in eq. (2), has a rank of at most two and is subject to a determinant-free constraint, meaning that a displacement-discontinuity source cannot exceed five DOFs. Indeed, the combined DOFs of a displacement discontinuity and an $n$-vector amount to five. Equation (2) restricts the direction of the opening mode to the fault-normal by imposing vanishing tangential-tangential components. Although a traceless condition is typically imposed to exclude tensile components when assuming pure shear slip, the tensile component by default serves to resolve the degeneracy between the fault normal and the displacement discontinuity within the moment/potency tensor.
Equation (2) is derived as an integrability condition, arising from the compatibility condition in continuum mechanics, for the surface distribution of inelastic strain (Appendix A). There is thus no physical requirement for it to be satisfied as long as the source is treated as inelastic strain within a finite volume. The excess degrees of freedom that cannot be represented by a surface capture the volumetric information. Therefore, eq. (1) should not be read as a mere restatement of the representation theorem for displacement discontinuities. It is instead a representation theorem for finite-volume sources where thickness is neglected. As demonstrated in the following section, the primary objective in geometry reconstruction from the potency is to extract the surface normal $n$ (eq. 3), defined in the displacement-discontinuity limit, from a source that is not strictly an ideal surface. For a point source, this task reduces to ignoring the eigenvalue with the smallest absolute magnitude to satisfy the determinant-free condition. For an areal source, it requires addressing the global constraints of eq. (3), which act through infinitesimal motions along the surface.
@@ 2.2 Decoupling surface reconstruction from potency estimation
The PDTI evaluates the potency density $D$ and the associated $n$-vector field through the well-known equations (1) and (2), and the $n$-vector field is translated into $\Gamma$ through eq. (3). Here is a problematic circular dependency between $D$ and $\Gamma$: the domain on which the potency density $D$ is defined is the fault surface $\Gamma$ (eq. 1), which is yet to be evaluated by $D$ (eq. 3). To infer $D$ prior to reconstructing $\Gamma$, the representation theorem (eq. 1) must be projected onto a fixed surface, here termed the reference surface, denoted by $\Gamma^{\circ}$ (Fig. 1).
@@ 3 Formulation of inverse mapping from potency density to fault shape
Equation (7) states that the cotangent of the $n$-vector corresponds to the downward gradient of a curve in a two-dimensional space, where the slip direction perpendicular to the $n$-vector is exactly the fault orientation. Accordingly, Shimizu et al. (2021) averaged the potency along the short axis of the reference surface and smoothly connected the resulting $n$-vector along the long axis to extract the two-dimensional features of a uniaxially bent fault (e.g., along-strike strike angle variations and along-dip dip angle variations). The same procedure adhering to eq. (6) can also reproduce quasi-two-dimensional shapes including uniaxial twists such as along-strike dip variations and along-dip strike variations (§5.3). This study regards such surface reconstruction fully governed by eq. (6), including quasi-two-dimensional surface reconstruction, as falling within the scope of the previous method by Shimizu et al. (2021).
@@
Then, eq. (13) requires the limit that places $\bf n$} in the integrable nearest neighbor of non-integrable $\bf n$$_{*}$, that is, the displacement discontinuity that best approximates given inelastic strain. It is termed dislocation limit hereafter. In eq. (14), the dislocation limit is described as
@@
When the non-integrable part $(\partial_{x}\partial_{y}-\partial_{y}\partial_{x})z$ expected from $n_{*}$ is small, the vector $n$ obtained from eq. (13) does not deviate significantly from $n_{*}$. Basically, we suppose this regime and proceed by adopting the prior $P(\Gamma|{\bf a})$ written as eq. (14). However, the overconstraint in surface reconstruction, where no smooth surface satisfies the forward model eq. (2) that leads to eq. (6), raises the question of what priors $P(\Gamma|{\bf a})$ are appropriate to relax eq. (2) when the non-integrability of $n_{*}$ is not small. It is important to note that the a priori constraint here is imposed on $\Gamma$ given $\bf a$}, rather than on the inelastic strain $\bf a$} distributed over a given surface $\Gamma$ as in slip inversion. Considering that $\bf a$} represents the $n$-vector, this approach of $\Gamma|{\bf a}$ implies a reversal of the standard definition of the $n$-vector, where $n$ is defined for $\Gamma(z)$. The PDTI is commonly formulated in such a way that determines $\Gamma$ from $\bf a$}; $\bf a$} is inferred by using another prior constraint independent of $\Gamma$. $P(\Gamma|{\bf a})$ introduced here was designed to align with this structure. While this approach reverses the standard definition of the $n$-vector, the physical nature of faulting supports this causal order of $\Gamma|{\bf a}$ rather than contradicts it. Even before the $n$-vector is disturbed by measurement errors that do not satisfy eq. (6), an earthquake source as a surface is an approximation of a fracture zone, an idealization following the nature of brittle failure localized within a quasi-face (Matsu’ura et al., 2019), so the exact observance of eq. (2) is unphysical from the outset. Thus, if read as a forward model that precedes the observation, $P(\Gamma|{\bf a})$ that determines a surface from a given $n$-vector field is a representation of the ambiguity in defining an earthquake fault surface. This surface is an approximation for a thin sheet of inelastic deformation.
@@
In summary, the potency density that best explains the data retains information that cannot be mapped back onto a displacement discontinuity, a limit of inelastic strain strictly localized on a surface. The resulting surface is forced to violate the integrability of the fault slope, which means the non-integrability of the potency density, the dislocation. Briefly, the best earthquake source is generally not a displacement discontinuity. In conventional point-source representations, the discrepancy between a six-degree-of-freedom moment tensor and a five-degree-of-freedom dislocation is resolved locally by simply discarding the eigenvalue with the smallest absolute magnitude. By contrast, in an areal finite-source representation, this dimensionality gap manifests as a non-local geometric conflict. Physically, when a normal vector $n$ is assigned to each coordinate based on local potency, the resulting field fails to satisfy the gradient-based constraints that couple adjacent points. The reconstructed source thus represents not a smooth, continuous fault, but a collection of infinitesimal fragments that are “torn everywhere”, a fragmented geometry where the dislocation is nowhere integrable (eq. 12). Therefore, reconstructing a meaningful fault geometry requires more than a local approximation, necessitating a global approach to the overdeterminacy inherent in the transition from an areal inelastic strain field to a displacement-discontinuity surface. Equation (11) represents the integrability condition necessary to project the areal inelastic strain onto displacement discontinuity without loss of information. It ensures the global consistency of the normal vector field, functioning as the areal-source counterpart to the determinant-free condition in point-source analysis.